OpenStudy (anonymous):

divide and check x^4+x^2 over x

OpenStudy (anonymous):

Factorize by x then you'll have: \[\frac{ x(x^3+x) }{ x } = \frac{ \cancel x (x^3+x) }{ \cancel x } = x^3+x\]

OpenStudy (anonymous):

thank you